### v1

# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing – original draft

Abstract
Flight delay prediction has become a key focus in air traffic management, as delays highlight inefficiencies that impact overall network performance. This paper presents a lightweight large language model-based multimodal flight delay prediction, formulated from the perspective of air traffic controllers monitoring aircraft delay after entering the terminal area. The approach integrates trajectory representations with textual aeronautical information, including flight information, weather reports, and aerodrome notices, by adapting trajectory data into the language modality to capture airspace conditions. Experimental results show that the model consistently achieves sub-minute prediction error by effectively leveraging contextual information related to the sources of delay. The framework demonstrates that linguistic understanding, when combined with cross-modality adaptation of trajectory information, enhances delay prediction. Moreover, the approach shows practicality and scalability for real-world operations, supporting real-time updates that refine predictions upon receiving new operational information.
keywords:
[…]
1 Introduction
The volume of air traffic grows in parallel with economic development. This growth can result in traffic density that exceeds the capacity of current air traffic management (ATM) systems, compromising both safety and operational efficiency [1, 2]. Such inefficiencies lead to flight delays that increase airline operating costs, burden airport coordination, disrupt passenger itineraries, and create network-wide ripple effects. The Federal Aviation Administration (FAA) defines delay as the excess time duration incurred during flight operations. These delays are divided into five operational phases, namely gate delay, taxi-out delay, en-route delay, terminal delay, and taxi-in delay [3]. Delays can also be classified by cause, including carrier delays, weather delays, security delays, and airspace-related delays [4]. The delays across operational phases and causes originate from different parts of the ATM system. Understanding relevant factors is essential for accurate, phase-specific delay prediction. This study focuses on delays after the terminal phase, which are of particular interest to air traffic controllers (ATCs), as monitoring and mitigating delays within this phase is their direct responsibility.
In ATM operations, the ATCs monitor trajectory data and textual aeronautical information, which belong to different modalities. Although these data are comprehensively human-interpretable, managing, inferring, and making decisions based on large volumes of such multi-modal information can still be overwhelming for ATCs. The growing public availability of large-scale aviation data has enabled the development of data-driven models to assist ATCs in these tasks, enhancing autonomy while preserving effective human–machine interaction. Textual information, such as flight schedules, weather reports, weather forecasts, and temporary operational notices, provides essential context for aviation operations and is nowadays accessible via many public databases. Aircraft trajectory data, which describes aircraft movement and operational behavior within controlled airspace, can be extracted from recorded Automatic Dependent Surveillance–Broadcast (ADS-B) databases. Integrating and understanding these contexts plays an important role in improving airspace-related delay prediction.
Since the operational context comprises information in different modalities, each requires distinct handling techniques, while their combined use requires effective data integration. For textual data, large language models (LLMs) have recently gained recognition as powerful models capable of understanding natural language. Pretrained on large-scale language corpora, LLMs can often interpret aviation-related text, such as flight descriptions and aeronautical information. On the other hand, trajectories are complex, high-dimensional time series that can be represented in a more compact form, which helps downstream models perform better even without further fine-tuning. However, trajectory data is not the type of input recent LLMs are trained to handle directly; therefore, integrating such data with LLMs requires additional adaptation techniques. This paper’s proposed method incorporates two pre-trained models, an LLM and a trajectory representation model, using a simple yet effective adaptation technique to integrate their embeddings. We formulate flight delay prediction as a multimodal regression problem, where both textual and trajectory data serve as inputs for prediction. The contributions of this paper are as follows:
  • We propose a language-trajectory multimodal framework for predicting delays beyond the terminal phase, incorporating comprehensive operational context through the integration of textual aeronautical information and aircraft trajectories, leveraging an LLM and a trajectory representation model for their respective modalities.
  • This paper uncovers that the linguistic understanding embedded in pretrained LLMs is crucial and can be effectively transferred and adapted to the flight delay prediction task, highlighting their potential for efficient multimodal inference without requiring computationally intensive full fine-tuning.
  • We demonstrate that incorporating a wide range of aeronautical information into the operational context leads to more accurate delay predictions, as the model can account for contextual factors that describe the situation in the airspace, which in turn extend the time an aircraft spends within the area.
  • Our framework supports real-time delay monitoring by processing continuously updated inputs, such as surveillance broadcast signals, enabling second-by-second updates that are useful for operators who require more frequent real-time updates as the aircraft approaches the airport.
2 Related Works
2.1 Delay Prediction
Flight delay prediction has long been an area of interest in the ATM community. Early studies employed tabular-based approaches, in which flight details (e.g., carrier, route, schedule, delay history) and contemporaneous weather data (e.g., visibility, wind, temperature) were extracted into structured tables and used as fixed-dimensional numerical or categorical features for machine learning models. Delay prediction can be formulated as a classification problem, where the goal is to determine whether a delay will occur. For example, prior studies have employed logistic regression [5] and a modified Random Forest model [6]. Meanwhile, [7, 8, 9, 10] have compared various non-neural machine learning models for delay classification. Integrating traffic count features, [11] predicts departure delays using linear regression, support vector machines (SVM), Multilayer Perceptron (MLP), and LightGBM. The delay prediction can also be formulated as a regression problem that predicts delay duration in minutes. For example, SVM [12] and an MLP using only flight information [13] have been applied to predict delay duration. [14] compares several models for predicting delay distributions from flight features; subsequently, [15, 16, 9] compare various models for delay regression.
Many works have incorporated historical data for delay classification in sequence. For instance, [17, 18] both employ long short-term memory (LSTM) models using flight information and weather data to classify next-day delay. However, the day-level granularity in both limits their real-time applicability. Complex network models, such as the LSTM-integrated random forest model [19] and the Convolutional Neural Network–Long Short-Term Memory (CNN-LSTM) [20], were employed to predict delays based on weather, delay propagation, and airport congestion, as measured by traffic count. [21] models multi-flight delay propagation using attention-based architectures that capture flight chain dependencies across sequential operations of the same aircraft, accounting for temporal coupling and propagation effects in aircraft rotations. [4] employs Gaussian Process Regression to predict flight delays using only the prior delays and flight dates, with the model learning the mean and covariance from historical delay patterns.
Several recent studies have adopted graph-based approaches for delay regression. [22] employs a GCN-CNN architecture, where a Graph Convolutional Network (GCN) captures spatial dependencies in a time-evolving network of airports, and a CNN models daily and weekly delay patterns. Similarly, [23] proposes a hybrid framework that combines diffusion GCNs with residual Gated Recurrent Units (GRUs), incorporating federated learning across airport networks for accurate, privacy-preserving forecasting using historical delay data. [24] represents air traffic as a directed multigraph for predicting delays caused by holding maneuvers. [25] employs an attention-based LSTM model combining aircraft trajectory with flight and weather data to classify delays as early, on-time, or late.
Table 1: Comparison of Prediction Contexts and Formulations in Delay Prediction Models
Method Prediction Context Formulation
Flight Info Weather NOTAMs Trajectory Data Airport Corr. Classif. Regr.
Current Forecast Focusing Active Prior
Tabular-Based Models
Linear Regression [15, 16] num num - - - - - - -
Logistic Regression [5, 7, 8, 10] num num - - - - - - -
Decision Tree [7, 10, 16] num num - - - - - -
Random Forest [7, 6, 8, 9, 10, 16, 14] num num - - - - - -
KNN [7] num num - - - - - - -
SVM [7, 12, 16] num num - - - - - -
Naïve Bayes [7, 8, 10] num num - - - - - - -
Gradient Boosting [7, 14, 8, 16, 9] num num - - - - - -
MLP [8, 9, 13, 14] num num - - - - - -
LR/SVM/MLP/LGBM [11] num num - - num* num* - - -
Sequential Models
LSTM [17, 18] seq seq - - - - - - -
ST-Random Forest [19] seq seq - - seq* seq* - seq -
CNN-LSTM [20] seq seq - - seq* seq* - seq -
CBAM [21] seq - - - - - - - -
Gaussian Process [4] seq - - - - - - - -
Graph-based Models
GCN-CNN [22] graph-seq - - - - - - graph-seq -
GCN-GRU [23] graph-seq - - - - - - graph-seq -
GATs [24] graph - - - - - - graph -
Trajectory-Based Models
LSTM-Attention [25] num num - - ts - - - -
Ours txt txt txt txt ts ts ts - -
Abbreviations: num = fixed-dimensional numerical/categorical tabular features, seq = sequence of numerical features, ts = raw time-series trajectory, txt = textual data, graph = graph-structured data, graph-seq = sequence of graph-structured data. *Only trajectory count, disregard full trajectory information.
2.2 Large Language Models
The Transformer architecture [26] has introduced powerful contextualization and has become the foundation for most subsequent language models. Pretrained language models began with GPT-1 [27, 28], which used generative pretraining with a causal decoder-only Transformer. BERT [29] introduced masked language modeling and next sentence prediction with a bidirectional encoder, while T5 [30] employed an encoder-decoder architecture for text-to-text tasks on the C4 corpus. The LLM was further advanced by GPT-3 [31], which demonstrated strong few-shot prompting without task-specific fine-tuning. This approach is now widely adopted in research and applications, along with techniques such as instruction tuning [32] for improved prompt-following and Reinforcement Learning from Human Feedback (RLHF) [33]. With these techniques, models such as GPT-3.5, GPT-4, GPT-4o, Gemini, Claude, and DeepSeek V3 were released and remain proprietary, serving as high-performing models for research and industry applications. Open-source LLMs have been released, enabling broader access within the communities, including Meta’s LLaMA series [34, 35, 36], Mistral-7B and Mixtral-8x7B [37, 38], Google’s Gemma series [39, 40, 41], Microsoft’s Phi models [42, 43, 44], the Qwen family [45, 46, 47, 48, 49], and DeepSeek-R1 [50]. These models achieve performance comparable to proprietary ones and unlock new possibilities for customization and specialized use through flexible adaptation.
2.3 Time Series and Multimodal Models
Given that trajectories are inherently time series, the development of time-series pretrained models, also referred to as time-series representation learning, can be traced back to the autoencoder [51], which is widely used in the ATM field. The encoder is trained to compress input sequences, and the decoder is trained to reconstruct the original data. Contrastive representation learning aligns similar samples in the embedding space while separating dissimilar ones. Notably, recent works include TS2Vec [52], which employs hierarchical contrastive learning between two randomly masked views; TF-C [53], which aligns time-based and frequency-based views; SimMTM [54], which utilizes masked time-series modeling; and InfoTS [55], which leverages meta-learning for data augmentation method selection. Aircraft Trajectory Segmentation-based Contrastive Coding (ATSCC) [56] is a framework for aircraft trajectory data that ensures consistent embedding of cumulative trajectory information within segments. It generates high-fidelity instance-level representations that outperform prior methods in instance-level tasks such as classification and clustering, with embeddings that align closely with human-defined aeronautical procedures.
LLMs and Large Multimodal Models (LMMs) for time series are rapidly growing research areas. Cross-modality adaptation is a powerful technique for leveraging pretrained models from one modality and adapting them for use in other modalities. LLM4TS [57] proposed a two-stage fine-tuning, where an LLM is first pretrained to align with the time-series modality, and then further fine-tuned on a forecasting task. LLMTIME [58] demonstrated a procedure for tokenizing continuous-valued time series into discrete tokens, enabling the direct use of pretrained LLMs for time series forecasting. Time-LLM [59] introduces a reprogramming layer that converts time series into LLM-friendly embeddings, combined with a prompt-as-prefix strategy and a trainable projection layer on top of a frozen LLM.
2.4 Large Language Model for Aviation Applications
LLM applications have extended to aviation-related language tasks. AviationGPT [60] fine-tuned and instruction-tuned LLMs on an aviation corpus comprising books, documents, and textual data such as METAR (Meteorological Aerodrome Reports), TAF (Terminal Aerodrome Forecasts), and NOTAM (Notice to Airmen). Aviation-BERT [61] was pretrained using MLM on reports from the Aviation Safety Reporting System and the National Transportation Safety Board. AeroBERT [62, 63] was introduced for classification and named entity recognition on engineering requirements. In [64], GPT-3.5-Turbo was fine-tuned to predict means of compliance for aerospace systems. CHATATC [65] was fine-tuned on Ground Delay Program data to be a conversational agent. LLMs have also been applied to various ATM tasks. In [66], GPT‑4o was fed with auto-generated flight planning prompts, wind hazard polygons, and operator preferences to generate route recommendations, which are then verified by the operator. The work in [67] presented supervised fine-tuning of LLMs, enabling the generation of conflict-free flight slot allocation. The ATM agents in [68] use a dual LLM-based framework that interacts with the BlueSky [69] simulator to monitor traffic, interpret situations, and generate conflict-minimized instructions. The application of LLMs for trajectory data has also been explored in recent works. FlightBERT [70] uses a bidirectional MTM to learn representations for flight trajectory prediction. FlightBERT++ [71] extends this to support additional downstream tasks, including anomaly detection and intent classification. The study in [72] fine-tunes the LLM on a text-reformatted prompt-based trajectory dataset. The work in [73] fine-tunes LLaMA to reconstruct aircraft positional states from incomplete sequences. SIA-FTP [74] integrates ATC speech and flight trajectory data by using BERT for text and FlightBERT++ for trajectories, fusing their representations for trajectory prediction.
3 Methodology
3.1 Problem Formulation
Figure 1: Flight Delay Accumulation of Aircraft Operations from Departure to Arrival
Considering a flight with identifier $i$, the time horizon of its entire operation is illustrated in Figure 1. Generally, the key temporal markers that define the operational timeline of a flight include the scheduled and actual departure times, airborne time, airspace entry time, landing time, and arrival time. As described in [3], delay can be decomposed into components corresponding to each phase of flight operations. As shown in Figure 1, the total delay is the accumulation of delays that occur during each phase of the flight. In this paper, we define the pre-terminal phase as all operations prior to entering the airspace. Accordingly, the pre-terminal delay consists of gate delay, taxi-out delay, and enroute delay. Gate delay may result from service-related operations, delay of previous flights, or other logistical factors. Taxi-out delay is commonly caused by ground holding procedures. Enroute delay is influenced by factors such as adverse enroute weather conditions and required detours. These components can be rearranged as follows:
$D_{pre}$ $=D_{gate}+D_{taxi-out}+D_{enroute}$ (1)
$D_{pre}$ $=D_{gate}+D_{taxi-out}+\left(T_{entry}^{act}-T_{entry}^{exp}-D_{gate}-D_{taxi-out}\right)$ (2)
$D_{pre}$ $=T_{entry}^{act}-T_{entry}^{exp}$ (3)
We define the post-terminal phase as the operations of aircraft after entering the airspace. The post-terminal delay encompasses terminal delay and taxi-in delay. Terminal delay typically arises from air traffic congestion, weather disruptions, or airspace restrictions. Taxi-in delay is generally attributed to ground congestion at the destination airport. These delay components can be rearranged as follows:
$D_{post}$ $=D_{terminal}+D_{taxi-in}$ (4)
$D_{post}$ $=\left(T_{land}^{act}-T_{land}^{exp}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)+\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{land}^{act}-T_{land}^{exp}\right)$ (5)
$D_{post}$ $=\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)$ (6)
Thus, the total delay is defined as the sum of delays accumulated during both the pre-terminal and post-terminal phases. This paper’s problem formulation is designed from the perspective of an ATC at the destination airport, who is responsible for monitoring aircraft within a designated control area. This involves tracking and updating delay information from the actual airspace entry time $T_{entry}^{act}$ until the actual arrival time $T_{arr}^{act}$. Since the scheduled arrival time is known, and $T_{entry}^{act}$ marks the start of monitoring, parts of the delay can be inferred from these known values.
$D_{total}$ $=D_{pre}+D_{post}$ (7)
$D_{total}$ $=\left(T_{entry}^{act}-T_{entry}^{exp}\right)+\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)$ (8)
$D_{total}$ $=T_{entry}^{act}-T_{arr}^{sch}+\left(T_{arr}^{act}-T_{entry}^{act}\right)$ (9)
$D_{total}$ $=T_{entry}^{act}-T_{arr}^{sch}+\Delta t_{post}$ (10)
The post-terminal duration is denoted as $\Delta t_{post}$. This value remains unknown during the monitoring period at the current observation time $t\in\left[T_{entry}^{act},T_{arr}^{act}\right)$ because $T_{arr}^{act}$ has not yet been observed. Thus, we aim to estimate $\Delta t_{post}$ using available contextual information. Unlike the graph-based and sequential models in Table 1, the past delays are accumulated in the gate delay for a particular flight, the component known to the ATC at the destination airport, and those terms are eliminated in our pre-terminal delay formulation in equation 3, leaving only the post-terminal duration ($\Delta t_{post}$) to be predicted. Accordingly, the estimated total delay is given by:
$\hat{D}_{total}=T_{entry}^{act}-T_{arr}^{sch}+\Delta\hat{t}_{post}$ (11)
The advantage of predicting the post-terminal duration is that it typically aligns with maneuvering procedures, as similar procedures usually result in similar durations; however, variations can still occur due to various factors. We hypothesize that accurate prediction of $\Delta\hat{t}_{post}$ requires the integration of flight information, weather conditions, non-periodic operational events, and traffic congestion. Therefore, comprising the most comprehensive context compared in Table 1, we define a scenario $s_{i,t}$ of the flight $i$ at time $t$ that consists of textual prompts that include general flight information, weather information, NOTAMs, and airspace conditions described by aircraft trajectories. Such that,
$s_{i,t}$ $=\left\{P^{fp}_{i,t},\ P^{wx}_{t},\ P^{no}_{t},\ X^{f}_{i,t},\ X^{a}_{i,t},\ X^{p}_{i,t}\right\}$ (12)
$P^{fp}_{i,t},\ P^{wx}_{t},\ P^{no}_{t}$ $\in\mathbb{Z}^{L^{(j)}_{i}},\quad j\in\{fp,wx,no\}$
$X^{f}_{i,t},\ X^{a}_{i,t},\ X^{p}_{i,t}$ $\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\text{max}}\times 3},\quad k\in\{f,a,p\}$
A scenario, $s_{i,t}$, comprises data from multiple modalities. The textual prompt $P^{fp}_{t}$ encodes general flight information, $P^{wx}_{t}$ captures weather conditions by combining METAR and TAF reports, and $P^{no}_{t}$ contains NOTAMs information. This paper introduces a trajectory-based characterization of airspace conditions, enabling the model to interpret congestion more effectively than the traffic count methods used in [11, 19, 20]. The trajectory data is represented as a time-series modality, where each timestep contains the $x,y,z$ coordinates in the ENU (East-North-Up) frame. To describe the airspace, we use three distinct groups of trajectories. The focusing trajectory $X^{f}_{i,t}$, the active trajectories $X^{a}_{i,t}$ those currently airborne at time $t$, and the completed prior trajectories $X^{p}_{i,t}$, which may provide additional context or relevant information. The prediction model is formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$, where $\hat{y}_{i}\equiv\Delta\hat{t}_{post,i}$, and the two notations are used interchangeably. The function $f_{\theta}$ represents the overall model, composed of several components. The text part includes a tokenizer $T(\cdot)$, followed by an embedding table $\text{Emb}[\cdot]$. The trajectory part consists of the ATSCC representation network, denoted as $f_{\text{atscc}}(\cdot)$ [56], followed by a trainable cross-modality adaptation network $f_{\text{xa}}(\cdot)$. The combined embeddings are then passed to a frozen LLM backbone $f_{\text{llm}}(\cdot)$. Finally, a trainable prediction head $f_{h}(\cdot)$ that predicts $\hat{y}_{i}$. The training of the model $f_{\theta}(\cdot)$ is formulated as a regression problem, minimizing the error between the predicted post-terminal duration $\hat{y}_{i}$ and the ground-truth $y_{i}$.
3.2 Dataset Preparation
The model training requires ground-truth pairs $(y_{i},s_{i,t})$, where $y_{i}$ denotes the actual post-terminal duration, and $s_{i,t}$ represents the multimodal input scenario. Monthly datasets are compiled for flights arriving at Incheon International Airport throughout 2022. To construct these datasets, we integrate information from multiple sources, including flight schedules, aircraft trajectories, meteorological data, and NOTAMs, resulting in a comprehensive multimodal collection for delay prediction.
3.2.1 General Flight Information Processing
We first construct a flight schedule database using data on arrival flights at Incheon International Airport for the year 2022, queried from Airportal [75]. The retrieved records consist of structured information, including scheduled arrival times, airline names, flight identifiers, and airport details such as departure and destination codes, as well as airport names. We matched the airports with their coordinates, including latitude, longitude, and altitude, and calculated the great-circle distance between the origin and destination. The flight haul type, categorized as short-haul, medium-haul, or long-haul, is determined based on this distance according to the classification provided in Wragg’s aviation dictionary [76]. The date and day of the week of arrival are recorded in Korea Standard Time. Certain fields are synthesized to complete the dataset. Aircraft type and registration are inferred based on the frequently used aircraft models on each specific route. The wake turbulence category is classified based on the aircraft type in accordance with ICAO’s Procedures for Air Navigation Services - Air Traffic Management Doc 4444 [77]. The actual airspace entry time and the current time stamp ($t$) are included here to provide additional context. Each flight $i$ is associated with general flight information, as detailed in Table 2, which is later used to construct the prompt $P^{fp}_{i,t}$ using the format in Figure 9.
Table 2: General Flight Information Data Descriptions and Data Types
Flight Information Description Data Type
airline_name_english Airline name in English string
callsign_code_iata IATA-format callsign string
callsign_code_icao ICAO-format callsign string
haul Flight haul type (short, medium, long, local) string
dep_code_iata Departure airport IATA code string
dep_code_icao Departure airport ICAO code string
dep_name_english Departure airport name string
dep_lat Departure airport latitude float64
dep_lon Departure airport longitude float64
dep_altitude Departure airport altitude (ft) float64
dest_code_iata Destination airport IATA code string
dest_code_icao Destination airport ICAO code string
dest_name_english Destination airport name string
dest_lat Destination airport latitude float64
dest_lon Destination airport longitude float64
dest_altitude Destination airport altitude (ft) float64
distance Great-circle distance between origin and destination (km) float64
actual_entry_time Time of first ADS-B appearance or airspace entry datetime64
sched_time_utc Scheduled arrival time (UTC) datetime64
date Scheduled Flight Date of arrival (KST) datetime64
day_of_week Scheduled Day of week of arrival (KST) string
aircraft_type Aircraft type string
aircraft_registration Aircraft registration code string
wake_turbulence_cat Wake turbulence category string
3.2.2 Textual Data Preprocessing
Since METARs, TAFs, and NOTAMs are provided in aeronautical coded text formats, we deliberately use their original, non-decoded forms to emphasize a key strength that the framework does not require separate decoders for these coded texts. Spanning the entire year of 2022, the METARs and TAFs were mass-downloaded from Ogimet [78], while NOTAMs were collected from the AIM Korea [79]. The active time periods of the coded texts vary by source. For a time $t$, the METAR report is retrieved by selecting the one whose release time falls within the preceding 30-minute update interval. In contrast, TAFs and NOTAMs explicitly define their active periods. TAF reports are typically issued every six hours, and multiple forecasts may overlap at a given time; therefore, we select the most recent valid TAF. For NOTAMs, multiple notices can be concurrently active, so we collect all active NOTAMs. The prepared METAR, TAF, and NOTAMs, at time $t$, are directly used to construct the prompts, $P^{wx}_{t}$ and $P^{no}_{t}$, as in Figure 10 and Figure 11.
3.2.3 Trajectory Data Preprocessing
The ADS-B data were sourced from the OpenSky database [80]. The data for both arrival and departure flights spanning the entire year of 2022 were queried using flight identification numbers from the Airportal website [75]. We extracted positional states from the surveillance data, including latitude, longitude, and barometric altitude. Barometric altitude was chosen for its lower noise levels and greater reliability compared to GPS vertical accuracy [81]; moreover, it is commonly used for traffic control. The positional features were transformed into ENU coordinates centered at the airport. The trajectories were resampled to 1-second intervals without interpolating into a fixed number of timestamps. Each trajectory was scaled by dividing by 120 kilometers, following the reproduction details in [56]. A preprocessed trajectory of flight $i$ consists of states $x_{i,t}=\{x_{i,t}^{x},x_{i,t}^{y},x_{i,t}^{z}\}$ over all recorded timestamps since its airspace entry. There are three types of trajectories: the focusing trajectory, active trajectories, and prior trajectories. Their temporal relationships are illustrated in Figure 2, and their definitions in a scenario $s_{i,t}$ are as follows:
  • Focusing trajectory: The trajectory of the flight $i$, comprising the states from its first ADS-B transmission in the airspace up to the current time $t$. It explains the current status and prior states of the focusing aircraft. The focusing trajectory is denoted as $X^{f}_{i,t}\in\mathbb{R}^{1\times\mathcal{T}^{f}\times 3}$, where $T^{f}$ is the number of trajectory states recorded from the aircraft’s first appearance in the airspace up to the current time $t$.
  • Active trajectories: The trajectories of other aircraft operating in the terminal maneuvering area (TMA) at time $t$, excluding the focusing flight $i$. This type of trajectory provides insight into the current airspace condition and congestion, similar to how an air traffic controller monitors multiple aircraft simultaneously. Each active trajectory contains states from the time the aircraft departed or first entered the TMA up to time $t$. The active trajectories are nan-padded and denoted as $X^{a}_{i,t}\in\mathbb{R}^{N^{a}\times\mathcal{T}^{a}_{\max}\times 3}$, where $N^{a}$ is the number of active trajectories at time $t$, and $\mathcal{T}^{a}_{\max}$ is the maximum trajectory length among all active trajectories.
  • Prior trajectories: Using the earliest timestamp among all active trajectories as a reference point, we collect all completed prior trajectories that were active at that time. Although these flights have completed their operations, they may still hold semantic importance. They capture patterns that may influence the paths of focusing and active flights, such as when aircraft are instructed to follow earlier ones. They also indirectly visualize ground traffic activity. The prior trajectories are nan-padded and represented as $X^{p}_{i,t}\in\mathbb{R}^{N^{p}\times\mathcal{T}^{p}_{\max}\times 3}$, where $N^{p}$ denotes the number of prior trajectories at time $t$, and $\mathcal{T}^{p}_{\max}$ is the longest trajectory length among them.
Therefore, the trajectories $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$ represent the query at time $t$ for flight $i$ under scenario $s_{i,t}$. Intuitively, these trajectories enable the model to achieve a comprehensive semantic understanding of the airspace, akin to the mental image used by human ATCs, without requiring explicit procedures such as trajectory classification or waypoint capacity estimation. This allows the model to directly perform the target task, estimating the post-terminal duration.
Figure 2: Temporal Relationship Between Focusing, Active, and Prior Trajectories
3.2.4 Label Preparation
The corresponding label $y_{i}$ for the regression task associated with $s_{i,t}$ was constructed as part of the supervised training setup. For flight $i$, the actual post-terminal duration, denoted as $\Delta t_{post,i}$, is calculated using the airspace entry time $T_{entry,i}^{act}$, derived from ADS-B data [80], taken as the moment the aircraft enters Incheon Airport’s TMA, and the actual arrival time $T_{arr,i}^{act}$, obtained from the official records in hour–minute format in Airportal [75]:
$y_{i}\equiv\Delta t_{post}=T_{arr}^{act}-T_{entry}^{act}.$ (13)
As a result, we obtain a multimodal dataset comprising general flight information in structured form, weather reports and NOTAMs in textual form, and trajectories as time-series data. Each component is integrated into $s_{i,t}$, which is then paired with a regression label $y_{i}$ representing the actual post-terminal duration. We repeat the preparation procedure for all flights $i$ within each month, selecting $\min(5,T_{i})$ time points per flight, where $T_{i}$ is the number of recorded states of flight $i$, to construct 12 datasets corresponding to each month spanning the year 2022. For each dataset, the most recent 10% are reserved as the testing set, another 10% is randomly sampled for validation, and the remaining scenarios are used for training. The visualization in Figure 3 illustrates the number of flights and the data size in each dataset.
Figure 3: Dataset Statistics
3.3 Neural Network Architecture
This section outlines the architecture of the proposed multimodal regression model $f_{\theta}(s_{i,t})$. This section describes how structured natural language prompts, derived from flight information, weather reports, and NOTAMs, as well as time-series trajectory data, are individually embedded. We also explain how a cross-modality adaptation network aligns trajectory and language embeddings, enabling seamless multimodal integration for predicting post-terminal duration through a regression head that replaces the LLM’s original language modeling head.
Figure 4: Overall Architecture of the Multimodal Prediction Framework
3.3.1 Natural Language Prompt Embedding
To process natural language inputs, including prompts for general flight information, weather information, and NOTAMs, we utilize the tokenizer and embedding table provided by the selected LLM. We denote the tokenizer as a function $T(\cdot)$, which transforms each combined prompt $P_{i,t}$, where $P_{i,t}=P^{\text{fp}}_{i,t}\,\|\,P^{\text{wx}}_{t}\,\|\,P^{\text{no}}_{t},$ into a sequence of subword tokens. The resulting tokenized sequence is $T(P_{i,t})=[p_{1},p_{2},\dots,p_{L}],$ where $L$ denotes the length of the token sequence. Each token $p_{l}\in T(P_{i,t})$ is then mapped to a continuous embedding vector $\mathbf{z}^{p}_{l}\in\mathbb{R}^{d}$ using the pretrained embedding table $\text{Emb}\in\mathbb{R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the embedding dimension. The output of this step is the embedded prompt representation:
$\mathbf{Z}^{p}_{i,t}=\text{Emb}[T(P_{i,t})]$ (14)
where $\mathbf{Z}^{p}_{i,t}\in\mathbb{R}^{L\times d}$ is the token-level embedding sequence for prompt $P_{i,t}$. The embedding table and tokenizer are kept fixed during training to preserve the pretrained language understanding, ensuring that natural language inputs are represented consistently and remain aligned with the pretrained LLM.
3.3.2 Trajectory Representation Model
The three trajectory types, focusing $X^{f}_{i,t}$, active $X^{a}_{i,t}$, and prior $X^{p}_{i,t}$, are time-series data compressed into instance-level representations using a pretrained trajectory encoder, where each embedding represents one trajectory. Thus, it is crucial that each embedding preserves the semantic fidelity of its corresponding trajectories and aligns well with human understanding of air traffic patterns. ATSCC [56] is well-suited for air traffic trajectory data, showing strong performance in instance-level tasks like classification and clustering, with representations that align with aeronautical procedures without relying on predefined labels. Therefore, the trajectory representation encoder model $f_{atscc}(\cdot)$ follows the architecture and was pretrained in accordance with the reproduction details of ATSCC.
We reproduce the trajectory encoder by strictly following the reproduction details of the ATSCC framework [56]. The causal transformer encoder architecture features 12 layers with a model dimension of 768, feed-forward network dimensions of 3072 using the Gaussian Error Linear Unit (GELU) activation, and 12 attention heads. The model has an input projection from 9 to the model dimension and an output projection from the model dimension to 320. ATSCC follows the NoPos [82], which omits positional encoding. Additionally, L2 normalization is applied after both input and output projections. The model is pre-trained on ATFMTraj [83] data, using 5-second interval downsampling and combining both Incheon departures and arrivals. To enhance the expressiveness of trajectory data, additional geometric features, including direction vectors and polar coordinates, are computed. For each timestep $k$ in a time-series instance, the additional geometric features are calculated as:
$\{x_{i,k}^{u_{x}},x_{i,k}^{u_{y}},x_{i,k}^{u_{z}}\}$ $=\frac{x_{i,k+1}-x_{i,k}}{\left\|x_{i,k+1}-x_{i,k}\right\|},$ (15)
$\{x_{i,k}^{r},x_{i,k}^{\sin\theta},x_{i,k}^{\cos\theta}\}$ $=\left\{\sqrt{\left(x_{i,k}^{x}\right)^{2}+\left(x_{i,k}^{y}\right)^{2}},\,\sin\left(\arctan 2(x_{i,k}^{y},x_{i,k}^{x})\right),\,\cos\left(\arctan 2(x_{i,k}^{y},x_{i,k}^{x})\right)\right\}.$ (16)
$x_{k}$ $=\{x_{i,k}^{x},x_{i,k}^{y},x_{i,k}^{z},x_{i,k}^{u_{x}},x_{i,k}^{u_{y}},x_{i,k}^{u_{z}},x_{i,k}^{r},x_{i,k}^{\sin\theta},x_{i,k}^{\cos\theta}\}$ (17)
During pretraining, we apply a dropout rate of 0.35 and binomial input embedding masking with a probability of 0.2. The model is trained using the AdamW optimizer with a batch size of 16, a learning rate of $1\times 10^{-5}$, and weight decay of $1\times 10^{-5}$. A grid search is conducted by separating a small portion of the dataset for validation to determine the optimal segmentation threshold and contrastive loss temperature based on classification and clustering performance. After pretraining, the trajectory encoder weights are frozen and not updated further.
Figure 5: Architecture of Causal Transformer Encoder Trained with ATSCC
The pre-trained encoder, denoted as $f_{\text{atscc}}(\cdot)$, first applies geometric extraction functions to $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 3}$ to expand each trajectory state into a 9-dimensional feature vector $X^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times\mathcal{T}^{(k)}_{\max}\times 9}$ with 5-second interval downsampling, followed by input projection and a forward pass through the causal transformer backbone. The final timestep output is then linearly projected to a 320-dimensional instance-level representation. The process of encoding the trajectory for type $k$ is illustrated in Figure 5 and formally defined as
$Z^{(k)}_{i,t}=f_{\text{atscc}}(X^{(k)}_{i,t})$, (18)
where $k\in\{f,a,p\}$ denotes the trajectory group and $Z^{(k)}_{i,t}=[z^{(k)}_{1},\dots,z^{(k)}_{N^{(k)}}]\in\mathbb{R}^{N^{(k)}\times 320}$ is the resulting representation. According to Figure 4, the trajectory representation encoder is placed at the beginning of the trajectory processing pipeline. Importantly, its weights are frozen throughout the training pipeline, meaning it does not participate in backpropagation. This design choice allows us to pre-encode the trajectories into their instance-level representations $Z^{(k)}_{i,t}$ prior to training. Doing so, we significantly improve memory efficiency, as we avoid training a highly memory-intensive nested transformer architecture, where long time-series sequences are first encoded into fixed-size vectors and then further embedded into a language-style sequence model. In contrast, pre-encoding the trajectories enables more efficient training without compromising the ability to incorporate rich temporal trajectory information.
3.3.3 Cross-modality Adaptation Network
To enable integration between language-based prompt embeddings and trajectory-based time-series representations, we employ a cross-modality adaptation network, denoted as $f_{\text{xa}}(\cdot)$, which adapts the trajectory inputs from the time-series modality into the language modality. This network is structured as a lightweight MLP, designed to project instance-level trajectory representations into the language model’s embedding space. Specifically, given the output from the pretrained trajectory encoder $Z^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times 320}$, the adaptation network transforms this vector into the LLM-friendly dimension $d$ as follows:
$\mathbf{Z}^{(k)}_{i,t}=f_{\text{xa}}(Z^{(k)}_{i,t})$, (19)
where $\mathbf{Z}^{(k)}_{i,t}\in\mathbb{R}^{N^{(k)}\times d}$ is the LLM-compatible embedding of trajectory group $k\in\{f,a,p\}$ (focusing, active, or prior). The MLP architecture, which projects the 320-dimensional trajectory embedding to the LLM’s model dimension of $f_{\text{xa}}(\cdot)$, comprises two linear layers with GELU activation functions and a dropout rate of 0.35 applied after each activation. The cross-modality adaptation network is simple and efficient, introducing only a small computational cost while effectively bridging the gap between trajectory encoders and LLMs. Its ability to integrate with the LLM without architectural changes highlights the practicality of cross-modal integration with minimal engineering effort. The embeddings $\mathbf{Z}^{(k)}_{i,t}$ are inserted into designated positions within the tokenized static trajectory prompts, forming a unified input sequence of language and trajectory information processed by the LLM backbone. To combine the trajectories with textual embeddings, we insert trajectory embeddings between guiding prompts (Figure 12) to help the model distinguish different trajectory groups. The trajectory input is formally defined as:
$\mathbf{Z}^{T}_{i,t}=\text{Concat}(\text{Emb}[T(P_{st,1})],\ \mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],\ \mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],\ \mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})]),$ (20)
This concatenated sequence $\mathbf{Z}^{T}_{i,t}\in\mathbb{R}^{(L_{st,1}+1+L_{st,2}+N^{a}+L_{st,3}+N^{p}+L_{st,4})\times d}$ represents the complete trajectory-informed input for the multimodal model, where $L_{st,k}$ denotes the length of the $k$-th static prompt segment after tokenization and embedding, 1, $N^{a}$, and $N^{p}$ are added to the embedding sequence length, as we use the instance-level representation of each trajectory as a single token.
3.3.4 Large Language Model Backbone
To leverage pretrained linguistic understanding, we retain the LLM backbone while replacing its language modeling head with a regression head, enabling prediction of continuous post-terminal duration values. In this paper, we consider LLMs with sizes under 1.5 billion parameters, as the task is formulated as a single-value regression problem and does not require large-scale generative capacity. These small LLMs can run efficiently on a single consumer-grade GPU, demonstrating their suitability for deployment by a wider range of users. The candidate backbone models include LLaMA3.2-1B, LLaMA3.2-1B-Instruct, Qwen2.5-1.5B, Qwen2.5-1.5B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, TinyLlama-v1.1, TinyLlama-1.1B-Chat-v1.0, and Pythia-1B. All model parameters and the embedding table are kept frozen as provided by the original sources to preserve their pretrained linguistic capabilities and ensure memory efficiency.
The LLM backbone processes multimodal inputs consisting of textual prompt embeddings, $\mathbf{Z}^{p}_{i,t}$ obtained via the corresponding tokenizer and embedding table, and trajectory-informed embeddings $\mathbf{Z}^{T}_{i,t}$ generated by the representation and cross-modality networks. Therefore, the input sequence fed to the LLM backbone is formed by concatenating these two components:
$\mathbf{Z}_{i,t}=\text{Concat}(\mathbf{Z}^{p}_{i,t},\mathbf{Z}^{T}_{i,t}),$ (21)
The concatenated sequence $\mathbf{Z}_{i,t}$ is then passed through the LLM backbone:
$\mathbf{h}_{i,t}=f_{\text{llm}}(\mathbf{Z}_{i,t})[-1],$ (22)
and due to the autoregressive nature of causal language models, we extract only the last hidden state as the representation that summarizes all preceding tokens: $\mathbf{h}_{i,t}\in\mathbb{R}^{d}$, which is then used by the output regression head.
3.3.5 Output Regression Head
For the final regression output, we attach a lightweight MLP output head to the model, which maps the final contextual representation $\mathbf{h}_{i,t}$ to a scalar prediction. This vector $\mathbf{h}_{i,t}\in\mathbb{R}^{d}$ is a multimodal encoding of the scenario $s_{i,t}$, capturing fused semantics from flight information, weather information, NOTAMs, and trajectory data. The regression head comprises three linear layers: the first maintains the LLM hidden size, the second halves the dimensionality, and the third outputs a scalar prediction. Each of the first two layers is followed by GELU activation and a 0.3 dropout rate. Formally, the model output is expressed as:
$\hat{y}_{i}=f_{h}(\mathbf{h}_{i,t}),$ (23)
where $f_{h}(\cdot)$ denotes the MLP-based output head. This head allows the pretrained LLM backbone to be repurposed for regression tasks, efficiently producing a prediction from the rich multimodal context encoded in $\mathbf{h}_{i,t}$.
3.4 Training Pipeline
For each scenario $s_{i,t}$, the textual components, general flight information, weather reports, and NOTAMs, are tokenized and embedded to form the prompt embedding $\mathbf{Z}^{p}_{i,t}$. Separately, the static prompts that describe the structure and semantics of the trajectory groups (focusing, active, and prior) are combined with their corresponding instance-level representations, which were pre-encoded by the frozen ATSCC model $f_{\text{atscc}}$ and projected via the trainable adapter $f_{\text{xa}}$ to form the trajectory-informed sequence $\mathbf{Z}^{T}_{i,t}$. Those embedding sequences are concatenated into a unified sequence $\mathbf{Z}_{i,t}$, which is then fed into the frozen LLM backbone $f_{\text{llm}}$. Since the LLM is a causal model, only the final hidden state $\mathbf{h}_{i,t}$ is extracted, representing a context-aware summary of the entire input sequence. This hidden state is passed to a trainable regression head $f_{h}$ to produce a scalar prediction $\hat{y}_{i}$, corresponding to the estimated time spent in the airspace. The ground-truth target $y_{i}$ is standardized prior to training. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-5}$, weight decay of $1\times 10^{-5}$, and a mini-batch size of 4. The model is trained using the Smooth L1 Loss:
$\mathcal{L}_{\text{smooth}}(\hat{y},y)=\begin{cases}0.5\cdot(\hat{y}-y)^{2},&\text{if }|\hat{y}-y|<1\\ |\hat{y}-y|-0.5,&\text{otherwise}\end{cases}$ (24)
which is robust to outliers, well-suited, and in some cases, prevents exploding gradients [84]. The parameters of the pretrained components, namely the embedding table Emb, LLM backbone $f_{\text{llm}}$, and the trajectory encoder $f_{\text{atscc}}$, are frozen. Only the cross-modality adaptation network $f_{\text{xa}}$ and the regression head $f_{h}$ are updated during training, as illustrated in Figure 4, ensuring memory efficiency by avoiding gradient allocation to the LLM and trajectory models. All training and evaluation are conducted using Python 3.11.8 and PyTorch 2.0.1 with CUDA Toolkit 11.8, running on an NVIDIA GeForce RTX 4090 GPU.
Algorithm 1 Multimodal Training Pipeline for Flight Delay Regression
1:Input: Scenario dataset $\mathcal{S}=\{(s_{i,t},y_{i})\}_{i=1}^{N}$, where $s_{i,t}=\{P^{fp}_{i,t},P^{wx}_{t},P^{no}_{t},X^{f}_{i,t},X^{a}_{i,t},X^{p}_{i,t}\}$
2:Output: Trained model parameter $\theta$
3:for $s_{i,t}$ in $\mathcal{S}$ do
4:  $Z^{(j)}_{i,t}=f_{\text{atscc}}(X^{(j)}_{i,t}),\quad j\in\{f,a,p\}$
5:end for
6:for each epoch do
7:  for each batch $\mathcal{B}\subset\mathcal{S}$ do
8:   for each scenario $(s_{i,t},y_{i})$ in $\mathcal{B}$ do
9:    Build prompt $P_{i,t}=P^{fp}_{i,t}\,\|\,P^{wx}_{t}\,\|\,P^{no}_{t}$
10:    $\mathbf{Z}^{p}_{i,t}=\mathrm{Emb}[T(P_{i,t})]$
11:    $\mathbf{Z}^{(j)}_{i,t}=f_{\text{xa}}(Z^{(j)}_{i,t}),\quad j\in\{f,a,p\}$
12:    $\mathbf{Z}^{T}_{i,t}=\text{Concat}(\text{Emb}[T(P_{st,1})],\ \mathbf{Z}^{f}_{i,t},\ \text{Emb}[T(P_{st,2})],\ \mathbf{Z}^{a}_{i,t},\ \text{Emb}[T(P_{st,3})],\ \mathbf{Z}^{p}_{i,t},\ \text{Emb}[T(P_{st,4})])$
13:    $\mathbf{Z}_{i,t}=\text{Concat}(\mathbf{Z}^{p}_{i,t},\mathbf{Z}^{T}_{i,t})$
14:    $\mathbf{h}_{i,t}=f_{\text{llm}}(\mathbf{Z}_{i,t})[-1]$
15:    $\hat{y}_{i}=f_{h}(\mathbf{h}_{i,t})$
16:   end for
17:   $\mathcal{L}=\frac{1}{|\mathcal{B}|}\sum_{i}\text{SmoothL1Loss}(\hat{y}_{i},y_{i})$
18:   Backpropagate and update trainable parameters in $f_{\text{xa}}(\cdot)$ and $f_{h}(\cdot)$
19:  end for
20:end for
4 Results and Discussion
This section begins by outlining the evaluation metrics used to assess the model after updating the parameters of the cross-modality adaptation network $f_{\text{xa}}(\cdot)$ and the regression head $f_{h}(\cdot)$. We then compare model performance across different LLM backbones to examine how linguistic capability affects prediction accuracy. Finally, we conduct a prompt structure analysis to evaluate the importance of each type of aeronautical information and its contribution to delay prediction accuracy.
4.1 Evaluation Metrics
To quantify the model’s predictive accuracy, we first evaluate its ability to predict the post-terminal duration $y_{i}\equiv\Delta t_{post,i}$ directly from the scenario input $s_{i,t}$, formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$. Ultimately, we assess the model’s effectiveness in estimating the total delay $D_{i}$, by incorporating the known actual entry time $T^{act}_{entry,i}$ and scheduled arrival time $T^{sch}_{arr,i}$, expressed as $\hat{D}_{i}=\left(T^{\text{act}}_{\text{entry},i}-T^{\text{sch}}_{\text{arr},i}\right)+\hat{y}_{i}$. The performance of both predictions is evaluated using a set of standard regression metrics. These metrics collectively capture various aspects of model performance, including the average magnitude of prediction errors, their squared severity, relative error scaling, and the proportion of variance explained by the model.
  • Mean Absolute Error (MAE) measures the average magnitude of errors in predictions, without considering their direction. The evaluation on total delay is given by:
    MAE $=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{D}_{i}-D_{i}\right|$ (25)
    $=\frac{1}{N}\sum_{i=1}^{N}\left|\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\right|$ (26)
    $=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$ (27)
    Therefore, the evaluation on the total delay is equivalent to evaluating the prediction of the post-terminal duration. MAE is easy to interpret and robust to outliers, as it represents the average magnitude of errors in absolute terms. However, it may not adequately reflect the impact of larger errors due to its linear penalty formulation.
  • Mean Squared Error (MSE) squares each error term, thus giving more weight to larger errors. The MSE evaluation on the delay is calculated by:
    MSE $=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{D}_{i}-D_{i}\right)^{2}$ (28)
    $=\frac{1}{N}\sum_{i=1}^{N}\left(\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\right)^{2}$ (29)
    $=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}$ (30)
    Thus, similar to MAE, the calculation of MSE for the total delay yields the same result as that for the post-terminal duration. MSE is particularly sensitive to outliers due to its quadratic error term, making it well-suited for scenarios where larger deviations from the actual values are especially critical.
  • Root Mean Squared Error (RMSE) is the square root of MSE:
    $\text{RMSE}=\sqrt{\text{MSE}}$ (31)
    RMSE combines the benefits of MSE’s sensitivity to large errors with a more intuitive scale for assessment.
  • Symmetric Mean Absolute Percentage Error (SMAPE) measures the relative prediction error as a percentage. The SMAPE for delay prediction and post-terminal duration prediction is given as follows:
    $\text{SMAPE}_{D}$ $=\frac{100\%}{N}\sum_{i=1}^{N}\frac{\left|\hat{D}_{i}-D_{i}\right|}{\left(\left|D_{i}\right|+\left|\hat{D}_{i}\right|\right)/2}$ (32)
    $\text{SMAPE}_{\Delta t}$ $=\frac{100\%}{N}\sum_{i=1}^{N}\frac{\left|\hat{y}_{i}-y_{i}\right|}{\left(\left|y_{i}\right|+\left|\hat{y}_{i}\right|\right)/2}$ (33)
    Due to the absolute terms in SMAPE’s denominator, we separately evaluate delay prediction ($\text{SMAPE}_{D}$) and post-terminal duration prediction ($\text{SMAPE}_{\Delta t}$). SMAPE is used instead of MAPE to reduce instability from zero ground truth values, common in on-time flights. While SMAPE can still be unstable when true or predicted values are near zero, it is generally more robust than MAPE.
  • Coefficient of Determination ($R^{2}$) quantifies the proportion of variance in the dependent variable explained by the model. The $R^{2}$ scores for both delay and post-terminal duration prediction are computed as follows:
    $R^{2}_{D}$ $=1-\frac{\sum_{i=1}^{N}(\hat{D}_{i}-D_{i})^{2}}{\sum_{i=1}^{N}(D_{i}-\bar{D})^{2}}$ (34)
    $R^{2}_{\Delta t}$ $=1-\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$ (35)
    Due to the presence of the mean in the denominator, the $R^{2}$ score for total delay ($R^{2}_{D}$) cannot be equivalent to that of the post-terminal duration ($R^{2}_{\Delta t}$); thus, they are evaluated separately. The $R^{2}$ score provides an intuitive measure of goodness-of-fit, indicating the proportion of variance explained by the model. While it can be misleading for small or imbalanced datasets, this is less of a concern given the scale and diversity of our data.
4.2 Model Performance
This section illustrates the performance of our framework when integrated with various LLM backbones, allowing for a direct comparison of how the varying linguistic capabilities of pre-trained LLMs with different architectures and training strategies impact prediction accuracy. Our evaluation protocol trains the overall model while replacing the LLM backbone with various candidates. For each backbone, we conduct experiments on 12 monthly datasets from 2022, training and testing separately on each month’s corresponding train-test split, visualized in Figure 3. Each scenario $s_{i,t}$ contains the full context, including flight plan, weather, NOTAMs, and trajectories representing airspace conditions. This full context is provided to the model to indirectly predict the total delay $D_{i}$ through the estimation of the post-terminal duration $\hat{y}_{i}=f_{\theta}(s_{i,t})$, using hyperparameters strictly following the reproduction details.
Table 3 presents a performance comparison of several lightweight LLMs on the flight delay prediction task. Among the models evaluated, LLaMA-3.2-1B-Instruct consistently outperforms others across all key metrics for many months. We hypothesize that its high-quality pretraining corpus contributes to its ability to generalize to aeronautical information. Its instruction-tuned training enhances the ability to follow structured prompts, and despite having relatively few parameters, it remains capable of capturing long context in post-terminal duration prediction.
Across the evaluated models, those trained primarily as English single-language models, unlike the Qwen-based model, tend to outperform multilingual models, suggesting that linguistic alignment with the structured English prompts in our framework provides an advantage. Interestingly, DeepSeek‐R1‐Distill‐Qwen‐1.5B, although distilled from the math‐specialized Qwen‐2.5‐1.5B‐Math model, performs better than its Qwen counterparts, suggesting that the distillation process may have introduced useful generalization capabilities. The TinyLLaMA variants demonstrate competitive performance, although they generally do not achieve the accuracy of LLaMA-3.2 models. In addition, the Pythia‐1B model achieves strong results, in some cases exceeding LLaMA-3.2-1B-Instruct, suggesting that it can be a suitable candidate under appropriate training conditions.
Although all models are provided with the full context and trained until the trajectory embeddings align with the tokens that enable the LLM to predict delays with minimal error, different LLM backbones still exhibit varying levels of performance. This reflects the importance of the models’ linguistic understanding in achieving accurate post-terminal duration prediction. Thus, it is not only the inclusion of the full context that matters, but also the model’s ability to effectively interpret that context in order to achieve the best prediction results.
In addition to assessing the comparative effectiveness of each model, we demonstrate that our proposed framework is compatible with arbitrary LLM backbones, as it consistently integrates trajectory representations with the language modality for accurate delay prediction across different models. This highlights the potential of the proposed cross-modality adaptation technique for application to larger and more capable models.
Our framework achieves high $R^{2}_{D}$ values due to our unique problem formulation, which leverages information already available to ATCs at the destination aerodrome, namely, the scheduled time of arrival and the actual time when the aircraft enters the TMA, marking the start of monitoring. This enables efficient prediction without the need for historical delay data or airport network information, unlike the sequential and graph-based models compared in Table 1. By narrowing the focus to a single airport, the framework aligns with the operational scope of ATCs and achieves significantly more accurate delay prediction, without requiring the ability to monitor many airports simultaneously.
From a practical perspective, our framework with several backbones maintains prediction error within one minute and can be considered deployable, as delays are typically recorded at the minute level [75]. While second-level accuracy may be desirable for certain operational needs, the framework already meets the current standard of minute-level records, and future access to higher-resolution data could provide opportunities for even greater precision.
Figure 6: Test MAE Across Months and Models
Table 3: Average Model Performance Comparison
[…]
LLaMA3.2-1B 0.9911 2.2794 1.5036 11.3009 0.99980 7.9400 0.9134
LLaMA3.2-1B-Instruct 0.9214 1.9965 1.4111 10.6924 0.99982 7.1308 0.9244
Qwen2.5-1.5B 1.1840 3.2651 1.7989 12.8570 0.99971 9.0699 0.8763
Qwen2.5-1.5B-Instruct 1.1970 3.4163 1.8335 13.0869 0.99969 9.1841 0.8705
DeepSeek-R1-Distill-1.5B 1.0469 2.5713 1.6000 11.7409 0.99977 7.8834 0.9026
TinyLlama-v1.1 1.0268 2.4719 1.5684 11.5998 0.99978 7.9976 0.9059
TinyLlama-1.1B-Chat-v1.0 0.9988 2.3257 1.5192 11.3519 0.99980 7.8986 0.9117
Pythia-1B 0.9551 2.1284 1.4545 11.0689 0.99981 7.4839 0.9192
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
4.3 Effectiveness of Context Inclusion
A key contribution of our proposed framework is the integration of a comprehensive set of information, providing a representation of the operational environment to the model, including general flight information, weather reports and forecasts, NOTAMs, and airspace conditions modeled through three types of trajectories set. These contextual components are of varying importance; therefore, to evaluate their contributions, we adopt an ablation-based approach in which each component is selectively removed from the multimodal scenario representation $s_{i,t}$, followed by retraining and re-evaluation under strictly identical hyperparameter settings. We used Meta’s LLaMA-3.2-1B-Instruct as the LLM backbone, for the results presented in Table 3 show that the model consistently outperforms other LLMs across the majority of datasets and evaluation metrics. This controlled procedure enables us to directly quantify the performance degradation resulting from the absence of each context type, thereby providing empirical evidence for its contribution to the overall predictive capability. Moreover, this analysis was conducted to support our design choice of using full context in the final configuration, demonstrating the value of leveraging multimodal aeronautical data for accurate prediction of $\Delta t_{post}$. The importances of context elements are discussed as follows:
  • Texual Data: Since the textual data provides the context for flight information, weather, and NOTAMs, removing them consistently degrades performance. This indicates that the LLM relies on textual context in addition to trajectory features for prediction. While the trajectory is directly proportional to the post-terminal duration, the extension of this duration is strongly influenced by the surrounding operational context.
  • General Flight Information: The exclusion of basic flight information leads to a clear drop in performance, underscoring the importance of general operation details of the flight. This finding aligns with prior works, which predominantly rely on general flight information as the primary input for delay prediction, as shown in Table 1.
  • Weather Information: The exclusion of weather information, including both METAR and TAF reports, leads to a decrease in model performance. This observation is consistent with many prior studies that weather data is a critical input for delay prediction (Table 1). It also aligns with the analysis in [3], which identifies weather as a major factor contributing to flight delays and emphasizes its necessity as context for accurate prediction.
    • METAR Information Removing METAR while keeping TAF makes the model depend solely on forecasts. Having at least TAF still enables the model to perform better than removing all weather information, since TAF, when recently issued and accurate, can also approximate current weather conditions. However, without METAR, which provides real-time weather observation, the model’s performance slightly drops.
    • TAF Information We observe that removing TAF while retaining METAR can, in some months, improve performance, likely due to the reduced burden of long contextual inputs. However, on average, the absence of TAF leads to a larger performance drop than removing METAR, indicating that forecast information still contributes valuable signals for delay prediction beyond real-time observations.
  • NOTAMs: Excluding NOTAMs shortens the contextual input, which can sometimes reduce the burden of long sequences. However, it also removes special operational constraints or disruptions, such as runway closures or equipment failures, which can introduce sources of delay [3]. As a result, performance is noticeably reduced overall. This supports our design choice of incorporating NOTAMs as input for post-terminal duration prediction, thereby improving delay estimation.
  • Trajectory Data: Trajectory data is shown to be the most influential factor for post-terminal duration prediction, as its exclusion results in a severe reduction in performance. This is expected, since trajectory data inherently reflect real-time airspace awareness that cannot be inferred from static flight or weather information alone. From a technical perspective, removing trajectory inputs also eliminates the learnable portion of the model’s context. This suggests that integrating trajectory data contributes to the predictive accuracy of our framework.
    • Focusing Trajectories We find that removing the focusing trajectory leads to a large performance degradation, often comparable to the removal of all trajectory inputs and consistently more severe than excluding other trajectory types. This finding is expected and aligns with [25], as the focusing trajectory embedding captures the specific maneuvering procedures being executed and the state of the current flight under monitoring, which are most directly related to its post-terminal duration.
    • Active Trajectories Excluding the active trajectories from the combined input removes the model’s awareness of the current airspace conditions, particularly the level of traffic congestion and the overall traffic situation at that time. Therefore, the model cannot account for the congestion factor that often extends the post-terminal duration, making accurate prediction more difficult.
    • Prior Trajectories Since prior trajectories provide patterns that may influence the active and focusing trajectories and indirectly reflect ground traffic, excluding them leaves the model unaware of historical flow patterns in both air and ground operations, thereby reducing predictive accuracy.
The experimental results indicate that our accurate predictions are enabled by offering a different formulation from prior works and by incorporating more contextual information within a single airspace, as compared in Table 1. While long contextual inputs can sometimes cause the model to struggle, removing certain elements, such as TAFs and NOTAMs, from many datasets may even improve performance. However, including the full context generally ensures consistent and reliable results, as the model can rely on that information when it is actually required. Since LLMs function as selective sequence encoders with the ability to disregard irrelevant tokens, it is preferable to provide the full context and allow the model to ignore unneeded tokens rather than risk excluding potentially useful information.
The contribution of textual components follows a similar pattern, though the trajectory components play a more critical role. This is partly because trajectory features are trainable and partly because the post-terminal duration is strongly influenced by prevailing traffic conditions. Moreover, durations in other phases of flight are implicitly accounted for through events that occur during the pre-terminal phases. The framework demonstrates robustness even when certain contextual information is unavailable, underscoring its suitability for operational use. Without such a model, air traffic controllers would need to assume an average additional more than 3 minutes for every flight [85], which could introduce unnecessary errors for non-delayed aircraft. In contrast, our model achieves an average error between 0.9214 and 2.7395 minutes, highlighting its accuracy and operational value.
Figure 7: Test MAE Across Months and Context Removal Settings
Table 4: Prompt Component Ablation Results using Meta-LLaMA-3.2-1B-Instruct
[…]
Full Context 0.9214 1.9965 1.4111 10.6924 0.99982 7.1308 0.9244
Without Textual Information 1.0103 2.3610 1.5331 11.4995 0.99979 7.8583 0.9105
   Without Flight Information 0.9738 2.2846 1.4955 11.1673 0.99979 7.8250 0.9142
   Without Weather Information 0.9898 2.3286 1.5153 11.3504 0.99979 7.6066 0.9119
    Without METAR 0.9549 2.1245 1.4534 11.0112 0.99981 7.3700 0.9197
    Without TAF 0.9757 2.2584 1.4855 11.0981 0.99981 7.6145 0.9131
   Without NOTAMs 0.9573 2.1403 1.4428 10.8942 0.99982 7.5266 0.9188
Without All Trajectories 2.7395 16.3961 4.0383 23.3048 0.99852 22.4923 0.3826
   Without Focusing Trajectory 1.5997 6.7335 2.4641 15.6197 0.99944 14.2710 0.7387
   Without Active Trajectories 1.0137 2.3863 1.5368 11.6211 0.99979 7.9227 0.9098
   Without Prior Trajectories 1.0528 2.6147 1.6005 11.8261 0.99977 8.3614 0.9001
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Figure 8: Demonstration of Second-by-Second Delay Updates. Left: Full trajectory; top-right: predicted delay versus actual delay; bottom-right: absolute error over time.
4.4 Demonstration for Second-by-Second Delay Update
From the perspective of an ATC, delay information should be updated in real time. Unlike prior works, where delay is typically predicted at a coarse time scale, constrained by the low update frequency of input contexts such as weather data (e.g., every 30 minutes), our framework can provide second-level monitoring because it incorporates surveillance broadcast signals, which are transmitted frequently and thus allow the context to be updated in real time. In this section, we demonstrate the real-time capability of our model by using a network trained on January data and testing it on a sample trajectory from February. For each second along this trajectory, we construct a textual prompt that includes general flight information, weather reports, and NOTAMs, and we retrieve the corresponding focusing, active, and prior trajectories of the flight at that time. This setting challenges the model with entirely unseen data and evaluates its ability to provide continuous predictions. The selected trajectory, along with the prediction results and absolute errors over the terminal phase time horizon, is visualized in Figure 8.
Although the model underestimates the delay in this example, we observe that the absolute error generally remains within one minute, with predictions adapting as the flight progresses. Moreover, the variation in error over time suggests that the model refines its predictions as newly available information updates the prediction context. Importantly, maintaining sub-minute error levels is practically useful for ATM, since delays are typically recorded at the minute level [75]; keeping the error below one minute is therefore considered deployable. Thus, the real-time prediction model from our framework can potentially be integrated into post-terminal ATM systems to assist ATCs in estimating the delay of aircraft within the terminal area.
5 Conclusion
This paper presents a flight delay prediction framework that indirectly estimates post-terminal delays by predicting post-terminal duration. Since flight schedules and airspace entry times are already known to ATCs at the destination airport, the framework is designed to support their operational needs by providing accurate and continuously updated delay information. It employs a cross-modality adaptation technique that aligns multimodal data with LLMs, adapting trajectory embeddings from a pre-trained ATSCC encoder into the language modality. The training process remains lightweight, leveraging the pretrained knowledge of both the trajectory encoder and the LLM, while only updating a cross-modality adaptation network and an output regression head for prediction. The framework enables accurate post-terminal duration prediction, with performance gains driven by the comprehensive inclusion of aeronautical information, especially trajectory data, which contribute the most after the cross-modality adaptation. These inputs allow the model to reference key factors influencing delays, while the linguistic capabilities of LLMs enable the interpretation of complex contextual information. By incorporating a wide range of context within the narrower airspace setting, our framework addresses several key limitations of prior works, such as the reliance on prior delay information, the requirement for structured inputs, and the restriction to coarse update timescales. In contrast, our approach focuses on the real-time situation in the airspace, incorporates prior delays through information already available to ATC, flexibly integrates all necessary context, and supports second-by-second updates, making it particularly well-suited for real-time post-terminal ATM operations. Looking forward, the framework can be extended to other phases of delay prediction by incorporating contextual factors relevant to those phases. Furthermore, its generality opens opportunities for broader ATM tasks, such as trajectory prediction and conflict detection. Future work may also explore the integration of higher-quality operational data and additional modalities, such as images or audio, to further advance multimodal models in aviation research in a practical and easily deployable manner.
Appendix A Prompt Structure
A.1 General Flight Information Prompt
From the general flight information, which includes multiple data types as described in Table 2, we convert the data into a format compatible with language models using 10 distinct prompt templates. Examples of these prompt formats are illustrated in Figure 9. These templates convert structured data into natural language, aligning better with LLM pretraining knowledge. Moreover, using multiple formats improves prompt generalization of the trainable components by training the model with varied expressions of the same content.
General Flight Information Prompt Format Example Scheduled for arrival at {sched_time_utc} UTC on {date} ({day_of_week}), flight {callsign_code_iata}/{callsign_code_icao} by {airline_name_english} was set to land at {dest_name_english} ({dest_code_iata}/{dest_code_icao}, lat: {dest_lat}, lon: {dest_lon}, alt: {des_altitude} ft). The aircraft originated from {dep_name_english} ({dep_code_iata}/{dep_code_icao}, lat: {dep_lat}, lon: {dep_lon}, alt: {dep_altitude} ft), and the total route spanned {distance} km. The aircraft was a {aircraft_type} with registration {aircraft_registration}, and it belonged to the {wake_turbulence_cat} wake turbulence category. It was expected to enter airspace or appear via ADS-B at {actual_entry_time}. This was a {haul}-haul flight.
General Flight Information Prompt Format Example On {date}, a {haul}-haul flight by {airline_name_english} was scheduled to arrive at {dest_name_english} ({dest_code_icao}/{dest_code_iata}) at {sched_time_utc} UTC. Flight ID: {callsign_code_icao}/{callsign_code_iata}. The aircraft departed from {dep_name_english} ({dep_code_icao}, lat: {dep_lat}, lon: {dep_lon}, alt: {dep_altitude} ft) and was destined for {dest_name_english}, located at lat: {dest_lat}, lon: {dest_lon}, alt: {dest_altitude} ft. The aircraft was a {aircraft_type} with registration {aircraft_registration}, and it belonged to the {wake_turbulence_cat} wake turbulence category. The aircraft was expected to enter monitored airspace around {actual_entry_time}. Day of week: {day_of_week}, distance: {distance} km.
Figure 9: Examples of Prompt Format for General Flight Information Description
A.2 Weather Prompt
The weather prompt $P^{\text{wx}}_{t}$ is constructed from the METAR and TAF data queried at time $t$ retrieved during dataset preparation. These coded weather reports are parsed and reformatted into natural language prompts following a predefined structure. An example of the formatted weather prompt is shown in Figure 10.
Weather Information Prompt Format Example METAR in effect: 202202272230 METAR RKSI 272230Z 14006KT 5000 BR NSC 01/M00 Q1023 NOSIG= TAF in effect: 202202271700 TAF RKSI 271700Z 2718/2824 21006KT CAVOK TN01/2722Z TX08/2806Z TN03/2820Z BECMG 2721/2722 14006KT 4000 BR NSC BECMG 2800/2801 21006KT 6000 NSW BECMG 2814/2815 14010KT BECMG 2817/2818 17006KT 3500 -RA FEW010 BKN025 OVC100 TEMPO 2822/2824 29006KT 0800 DZ FG BKN005
Figure 10: Example of Formatted Weather Information Prompt
A.3 NOTAM Prompt
The NOTAM prompt $P^{no}_{t}$ is constructed from all active NOTAMs at time $t$, capturing time-specific operational constraints. Instead of decoding aeronautical codes, we maintain the raw text and format it using a fixed template (Figure 11), similar to the weather information prompt.
Special Notices Prompt Format Example Active NOTAMs: REF AIP SUP 31/22, A-CDM PHASE 2 TRIAL OPERATION PROGRESSIVE STEPS 2 AND 3 PERIOD CHG AS FLW 1. STEP 2 : FROM 1600 UTC 29 JUN 2022 TO 1600 UTC 11 JAN 2023 2. STEP 3 : FROM 1600 UTC 11 JAN 2023 TO 1600 UTC 27 JUN 2023) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372806.54N 1262700.83E, 372740.28N 1262724.18E 2. RADIUS : 10M 3. HGT : 20.24M(66.4FT) AMSL) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372812.81N 1262551.43E, 372815.70N 1262553.51E, 372814.91N 1262557.46E, 372812.15N 1262552.41E 2. RADIUS : 40M 3. HGT : 60.39M(198.13FT) AMSL) TDWR(TERMINAL DOPPLER WEATHER RADAR) U/S DUE TO WIP)
Figure 11: Example of Formatted Special Aerodrome Notices (NOTAMs) Prompt
A.4 Static Guiding Prompt for Trajectories
Three types of trajectories, $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$, represent distinct aspects of the airspace situation surrounding the flight of interest. To help the model distinguish and reason over these inputs, we introduce a structured prompt that explicitly labels each trajectory group. The guiding prompt segments specify which input embeddings correspond to each trajectory type, thereby helping the model distinguish their distinct semantic roles.
Trajectory Prompt $P_{st,1}$: Focus Trajectory Airspace is described using three trajectory types. This embedding is for the focus trajectory: {
Trajectory Prompt $P_{st,2}$: Active Trajectories } These embeddings are for other active trajectories: {
Trajectory Prompt $P_{st,3}$: Prior Trajectories } These embeddings are for past or inactive trajectories that may still matter:{
Trajectory Prompt $P_{st,4}$: Closing Prompt } Missing types will have no embedding. Based on the above, predict the total time spent in the airspace.
Figure 12: Static Prompt Segments for Trajectory Type Descriptions within Scenario Prompts
The static prompts (Figure 12) will be tokenized and embedded in parallel with the trajectory embeddings obtained through a trajectory representation model and a cross-modality adaptation network. The instance-level representations of the trajectories are inserted between the static prompt segments, resulting in a unified language-time series embedding that describes the airspace. This design mirrors how human operators visualize dynamic traffic, thereby enabling effective modeling of airspace conditions without dependence on tasks such as traffic count, traffic classification, capacity assessment, or conflict detection.
Appendix B Full Model Comparison Results
Table 5: Model Performance Metrics for Q1-Q2 (January–June)
[…]
January LLaMA3.2-1B 1.0037 2.4014 1.5497 8.4650 0.99990 8.3901 0.9048
LLaMA3.2-1B-Instruct 0.9074 2.1798 1.4764 7.7226 0.99991 7.5289 0.9136
Qwen2.5-1.5B 1.1204 3.0157 1.7366 9.3075 0.99987 9.0874 0.8805
Qwen2.5-1.5B-Instruct 1.1903 3.2394 1.7998 10.0064 0.99986 9.9379 0.8716
DeepSeek-R1-Distill-1.5B 1.0579 2.7419 1.6559 8.8657 0.99988 8.4753 0.8913
TinyLlama-v1.1 1.1592 3.2552 1.8042 9.4798 0.99986 9.7450 0.8710
TinyLlama-1.1B-Chat-v1.0 1.0482 2.6411 1.6251 8.9892 0.99989 8.2527 0.8953
Pythia-1B 0.9884 2.5769 1.6053 8.4581 0.99989 8.3523 0.8979
February LLaMA3.2-1B 1.0218 2.0502 1.4318 9.7268 0.99988 10.7475 0.9249
LLaMA3.2-1B-Instruct 0.9023 1.7515 1.3234 8.6933 0.99990 8.3369 0.9358
Qwen2.5-1.5B 1.2429 2.8620 1.6917 11.4294 0.99983 10.7699 0.8952
Qwen2.5-1.5B-Instruct 1.0868 2.3503 1.5331 10.4437 0.99986 9.5309 0.9139
DeepSeek-R1-Distill-1.5B 1.0078 2.1312 1.4599 9.7570 0.99987 8.8340 0.9219
TinyLlama-v1.1 1.0731 2.3178 1.5224 10.4845 0.99986 9.6429 0.9151
TinyLlama-1.1B-Chat-v1.0 0.9382 1.7698 1.3303 9.2389 0.99990 9.1174 0.9352
Pythia-1B 0.8619 1.5463 1.2435 8.6889 0.99991 8.0699 0.9434
March LLaMA3.2-1B 0.9014 1.5692 1.2527 8.6175 0.99989 7.3750 0.9410
LLaMA3.2-1B-Instruct 0.9054 1.9429 1.3939 8.3270 0.99986 7.5979 0.9269
Qwen2.5-1.5B 1.1308 2.9306 1.7119 10.0066 0.99979 9.0807 0.8898
Qwen2.5-1.5B-Instruct 1.0091 2.3622 1.5369 9.2749 0.99983 8.3688 0.9112
DeepSeek-R1-Distill-1.5B 1.0120 2.2487 1.4996 9.1955 0.99984 7.9032 0.9154
TinyLlama-v1.1 1.0384 2.3429 1.5306 9.4284 0.99983 8.1317 0.9119
TinyLlama-1.1B-Chat-v1.0 0.9774 2.2102 1.4867 8.8269 0.99984 7.7688 0.9169
Pythia-1B 0.9966 2.0688 1.4383 9.3237 0.99985 8.3512 0.9222
April LLaMA3.2-1B 1.1285 3.0418 1.7441 12.2291 0.99981 7.7010 0.8594
LLaMA3.2-1B-Instruct 0.9483 2.2204 1.4901 10.9033 0.99986 6.7337 0.8974
Qwen2.5-1.5B 1.1678 3.2390 1.7997 12.4321 0.99980 8.7932 0.8503
Qwen2.5-1.5B-Instruct 1.4183 4.1012 2.0251 14.5600 0.99975 9.9183 0.8105
DeepSeek-R1-Distill-1.5B 1.1252 2.9771 1.7254 12.1800 0.99982 7.8894 0.8624
TinyLlama-v1.1 1.1138 3.0726 1.7529 12.0094 0.99981 7.8906 0.8580
TinyLlama-1.1B-Chat-v1.0 1.0918 2.8578 1.6905 11.7179 0.99982 8.1742 0.8679
Pythia-1B 0.9781 2.4767 1.5738 11.1960 0.99985 6.7553 0.8856
May LLaMA3.2-1B 0.9708 2.3205 1.5233 10.3333 0.99983 7.2112 0.9155
LLaMA3.2-1B-Instruct 0.9294 2.1616 1.4702 9.9291 0.99984 6.5563 0.9213
Qwen2.5-1.5B 1.0664 2.8976 1.7022 10.9505 0.99979 7.4380 0.8945
Qwen2.5-1.5B-Instruct 1.0979 3.0247 1.7392 11.2851 0.99978 7.6943 0.8899
DeepSeek-R1-Distill-1.5B 1.0504 2.8495 1.6880 10.9981 0.99979 7.5171 0.8963
TinyLlama-v1.1 0.9860 2.4176 1.5549 10.3540 0.99982 6.9338 0.9120
TinyLlama-1.1B-Chat-v1.0 1.0561 2.6750 1.6356 11.0431 0.99980 7.6696 0.9026
Pythia-1B 0.9790 2.2545 1.5015 10.5039 0.99983 7.2104 0.9179
June LLaMA3.2-1B 1.0609 2.9319 1.7123 13.2001 0.99978 7.7398 0.8965
LLaMA3.2-1B-Instruct 0.9669 2.2254 1.4918 12.2753 0.99983 7.0361 0.9214
Qwen2.5-1.5B 1.1983 3.6029 1.8981 14.3700 0.99973 8.9778 0.8728
Qwen2.5-1.5B-Instruct 1.4008 4.6358 2.1531 15.9640 0.99965 10.2455 0.8363
DeepSeek-R1-Distill-1.5B 1.0752 3.0417 1.7440 13.1595 0.99977 7.3740 0.8926
TinyLlama-v1.1 1.0048 2.3494 1.5328 12.5272 0.99982 7.1316 0.9170
TinyLlama-1.1B-Chat-v1.0 1.0800 2.8636 1.6922 13.3478 0.99978 7.7455 0.8989
Pythia-1B 0.9670 2.1730 1.4741 12.4024 0.99984 6.9883 0.9233
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Table 6: Model Performance Metrics for Q3-Q4 (July–December)
[…]
July LLaMA3.2-1B 1.0013 2.3414 1.5302 11.9827 0.99981 7.5753 0.9056
LLaMA3.2-1B-Instruct 0.9779 2.0390 1.4279 11.8794 0.99983 7.3058 0.9178
Qwen2.5-1.5B 1.2220 3.7145 1.9273 13.7401 0.99970 8.6895 0.8502
Qwen2.5-1.5B-Instruct 1.0961 2.9431 1.7156 13.2183 0.99976 7.9123 0.8813
DeepSeek-R1-Distill-1.5B 1.0931 2.5127 1.5851 13.2139 0.99980 7.8593 0.8986
TinyLlama-v1.1 1.0093 2.2665 1.5055 12.3198 0.99982 7.5489 0.9086
TinyLlama-1.1B-Chat-v1.0 1.0159 2.1691 1.4728 12.1190 0.99982 8.0089 0.9125
Pythia-1B 0.9868 2.1418 1.4635 12.1528 0.99983 7.5604 0.9136
August LLaMA3.2-1B 1.0349 2.5877 1.6086 12.5538 0.99976 7.2465 0.9071
LLaMA3.2-1B-Instruct 0.9675 2.2742 1.5080 12.1986 0.99979 6.5976 0.9183
Qwen2.5-1.5B 1.2848 3.9831 1.9958 14.7317 0.99963 9.8714 0.8570
Qwen2.5-1.5B-Instruct 1.3044 4.5557 2.1344 15.0162 0.99958 8.9693 0.8364
DeepSeek-R1-Distill-1.5B 1.0830 3.1092 1.7633 12.8727 0.99971 7.5807 0.8884
TinyLlama-v1.1 1.0350 2.6723 1.6347 12.6425 0.99975 7.1527 0.9040
TinyLlama-1.1B-Chat-v1.0 1.0400 2.8707 1.6943 12.8569 0.99973 7.3788 0.8969
Pythia-1B 1.0538 2.6711 1.6343 13.0296 0.99975 7.3680 0.9041
September LLaMA3.2-1B 0.9571 2.1990 1.4829 11.0576 0.99979 8.4002 0.9224
LLaMA3.2-1B-Instruct 0.8721 1.8700 1.3675 10.3723 0.99982 6.8206 0.9340
Qwen2.5-1.5B 1.0753 2.8478 1.6875 12.1321 0.99973 8.6854 0.8995
Qwen2.5-1.5B-Instruct 1.0798 2.6380 1.6242 11.9620 0.99975 9.2549 0.9069
DeepSeek-R1-Distill-1.5B 0.9814 2.3100 1.5199 11.3042 0.99978 7.7500 0.9185
TinyLlama-v1.1 0.9202 2.1347 1.4610 10.7115 0.99980 7.6250 0.9247
TinyLlama-1.1B-Chat-v1.0 0.9475 2.0498 1.4317 10.8078 0.99981 8.0965 0.9277
Pythia-1B 0.9322 2.1866 1.4787 10.6819 0.99979 7.8955 0.9229
October LLaMA3.2-1B 0.9448 2.0228 1.4222 12.5445 0.99967 8.4875 0.9254
LLaMA3.2-1B-Instruct 0.8931 1.8767 1.3699 11.5245 0.99969 7.1497 0.9308
Qwen2.5-1.5B 1.0271 2.7741 1.6656 12.6564 0.99954 8.1672 0.8977
Qwen2.5-1.5B-Instruct 1.2379 4.2722 2.0669 14.6202 0.99929 10.0358 0.8424
DeepSeek-R1-Distill-1.5B 1.0357 2.5105 1.5844 12.9461 0.99959 8.3069 0.9074
TinyLlama-v1.1 0.9607 2.2818 1.5106 12.4053 0.99962 8.3639 0.9159
TinyLlama-1.1B-Chat-v1.0 0.9608 2.0309 1.4251 11.8308 0.99966 7.9350 0.9251
Pythia-1B 0.9059 1.8919 1.3755 11.5062 0.99969 7.3428 0.9302
November LLaMA3.2-1B 0.9526 2.1269 1.4584 12.1783 0.99965 7.3662 0.9252
LLaMA3.2-1B-Instruct 0.8932 1.7821 1.3350 11.4918 0.99970 7.4177 0.9373
Qwen2.5-1.5B 1.0650 2.4397 1.5620 13.3018 0.99959 8.1777 0.9141
Qwen2.5-1.5B-Instruct 1.0648 2.5090 1.5840 13.3741 0.99958 8.5437 0.9117
DeepSeek-R1-Distill-1.5B 1.0128 2.2492 1.4997 12.6010 0.99963 7.9037 0.9209
TinyLlama-v1.1 0.9379 1.9974 1.4133 12.1604 0.99967 7.5292 0.9297
TinyLlama-1.1B-Chat-v1.0 0.9117 1.8832 1.3723 11.8652 0.99969 7.8927 0.9338
Pythia-1B 0.8835 1.7301 1.3153 11.7209 0.99971 7.1682 0.9391
December LLaMA3.2-1B 0.9156 1.7595 1.3265 12.7216 0.99982 7.0403 0.9330
LLaMA3.2-1B-Instruct 0.8934 1.6342 1.2784 12.9921 0.99983 6.4888 0.9378
Qwen2.5-1.5B 1.6067 4.8744 2.2078 19.2261 0.99950 11.1006 0.8145
Qwen2.5-1.5B-Instruct 1.3777 4.3643 2.0891 17.3179 0.99955 9.7979 0.8339
DeepSeek-R1-Distill-1.5B 1.0280 2.1739 1.4744 13.7969 0.99978 7.2068 0.9173
TinyLlama-v1.1 1.0832 2.5549 1.5984 14.6752 0.99974 8.2762 0.9028
TinyLlama-1.1B-Chat-v1.0 0.9577 1.8877 1.3739 13.5796 0.99981 6.7434 0.9282
Pythia-1B 0.9285 1.8230 1.3502 13.1617 0.99981 6.7443 0.9306
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Appendix C Full Context Analysis Results
Table 7: Model Performance for Q1 (January–March)
[…]
January Full Context 0.9074 2.1798 1.4764 7.7226 0.99991 7.5289 0.9136
Without Textual Information 1.0977 2.8337 1.6834 9.2716 0.99988 9.6124 0.8877
Without Flight Information 0.9049 2.0748 1.4404 7.8173 0.99991 8.3820 0.9178
Without Weather Information 1.1516 3.0966 1.7597 9.5516 0.99987 9.7125 0.8773
Without METAR 0.9109 2.0230 1.4223 7.7339 0.99991 7.7235 0.9198
Without TAF 0.8787 1.9071 1.3810 7.3968 0.99992 7.7284 0.9244
Without NOTAMs 1.3565 3.9314 1.9828 10.8599 0.99983 11.0969 0.8442
Without All Trajectories 2.4455 12.7980 3.5774 16.7755 0.99946 22.4854 0.4927
Without Focusing Trajectory 1.1938 3.1098 1.7635 9.8255 0.99987 12.2782 0.8767
Without Active Trajectories 0.9296 2.2951 1.5150 8.2534 0.99990 8.1610 0.9090
Without Prior Trajectories 1.2892 3.7040 1.9246 10.9643 0.99984 10.9403 0.8532
February Full Context 0.9023 1.7515 1.3234 8.6933 0.99990 8.3369 0.9358
Without Textual Information 0.9457 1.8088 1.3449 9.3358 0.99989 8.6091 0.9337
Without Flight Information 0.8054 1.5574 1.2480 8.1958 0.99991 7.6528 0.9429
Without Weather Information 0.8104 1.3630 1.1675 8.0553 0.99992 7.3063 0.9501
Without METAR 1.0812 2.3881 1.5454 10.2895 0.99986 9.9683 0.9125
Without TAF 0.8408 1.4357 1.1982 8.1196 0.99992 7.8462 0.9474
Without NOTAMs 0.7789 1.2568 1.1211 7.9574 0.99993 6.6601 0.9540
Without All Trajectories 3.1719 19.9589 4.4675 22.8117 0.99883 29.1812 0.2688
Without Focusing Trajectory 1.4247 4.1994 2.0492 12.3865 0.99975 15.9053 0.8462
Without Active Trajectories 1.1284 2.4138 1.5537 10.8637 0.99986 9.4770 0.9116
Without Prior Trajectories 1.1932 2.7324 1.6530 10.7404 0.99984 10.6528 0.8999
March Full Context 0.9054 1.9429 1.3939 8.3270 0.99986 7.5979 0.9269
Without Textual Information 1.0915 2.8445 1.6866 9.8405 0.99980 8.9635 0.8930
Without Flight Information 0.9191 1.9582 1.3994 8.3935 0.99986 7.2947 0.9264
Without Weather Information 0.9575 2.1133 1.4537 9.0038 0.99985 7.5829 0.9205
Without METAR 0.9539 2.1857 1.4784 8.8464 0.99984 7.3382 0.9178
Without TAF 1.0184 2.3494 1.5328 9.2595 0.99983 8.2568 0.9116
Without NOTAMs 1.2039 3.1883 1.7856 10.4703 0.99977 9.5013 0.8801
Without All Trajectories 2.7423 15.4991 3.9369 19.6905 0.99889 23.4856 0.4170
Without Focusing Trajectory 2.8069 19.7132 4.4400 20.7654 0.99859 23.1440 0.2585
Without Active Trajectories 1.0162 2.3732 1.5405 9.4628 0.99983 8.6576 0.9107
Without Prior Trajectories 1.0102 2.2177 1.4892 9.3372 0.99984 9.3373 0.9166
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Table 8: Model Performance for Q2 (April–June)
[…]
April Full Context 0.9483 2.2204 1.4901 10.9033 0.99986 6.7337 0.8974
Without Textual Information 1.0218 2.6042 1.6138 11.6386 0.99984 7.2961 0.8797
Without Flight Information 0.9289 2.2644 1.5048 10.9751 0.99986 6.5958 0.8954
Without Weather Information 0.9853 2.4929 1.5789 11.2667 0.99985 6.8435 0.8848
Without METAR 0.9639 2.3067 1.5188 11.2155 0.99986 6.7523 0.8934
Without TAF 1.4467 4.4181 2.1019 14.6432 0.99973 10.9560 0.7959
Without NOTAMs 0.8963 2.0808 1.4425 10.4439 0.99987 6.4001 0.9038
Without All Trajectories 2.4288 13.6587 3.6958 21.3268 0.99916 18.2874 0.3689
Without Focusing Trajectory 2.4582 15.5982 3.9495 21.2149 0.99904 18.3237 0.2793
Without Active Trajectories 0.9459 2.2794 1.5098 10.9661 0.99986 6.6560 0.8947
Without Prior Trajectories 1.3311 3.7610 1.9393 14.0949 0.99977 9.6272 0.8262
May Full Context 0.9294 2.1616 1.4702 9.9291 0.99984 6.5563 0.9213
Without Textual Information 1.0176 2.3380 1.5290 10.6221 0.99983 7.1981 0.9149
Without Flight Information 0.9368 2.1732 1.4742 10.1804 0.99984 7.1578 0.9209
Without Weather Information 0.9742 2.1549 1.4680 10.2734 0.99984 7.5117 0.9215
Without METAR 0.9574 2.0493 1.4316 10.0460 0.99985 7.4294 0.9254
Without TAF 1.0062 2.4767 1.5738 10.7096 0.99982 7.2864 0.9098
Without NOTAMs 0.9029 1.9877 1.4099 9.4669 0.99985 6.9338 0.9276
Without All Trajectories 3.0113 17.8540 4.2254 24.0631 0.99868 23.2055 0.3499
Without Focusing Trajectory 1.7352 6.7964 2.6070 15.9799 0.99950 14.2820 0.7525
Without Active Trajectories 0.9798 2.0669 1.4377 10.4690 0.99985 7.2774 0.9247
Without Prior Trajectories 0.9710 2.3356 1.5283 10.1317 0.99983 7.3870 0.9150
June Full Context 0.9669 2.2254 1.4918 12.2753 0.99983 7.0361 0.9214
Without Textual Information 0.9924 2.3334 1.5275 12.4934 0.99982 6.9779 0.9176
Without Flight Information 1.2905 3.8926 1.9730 14.9703 0.99971 8.9481 0.8625
Without Weather Information 1.1042 3.1180 1.7658 13.6114 0.99976 7.3959 0.8899
Without METAR 1.0413 2.6701 1.6340 13.0183 0.99980 7.2847 0.9057
Without TAF 1.0177 2.8125 1.6771 12.3968 0.99979 7.8702 0.9007
Without NOTAMs 0.9211 1.8717 1.3681 11.7378 0.99986 6.4112 0.9339
Without All Trajectories 2.9809 18.3495 4.2836 25.6296 0.99861 21.9901 0.3520
Without Focusing Trajectory 1.3339 4.1165 2.0289 14.6952 0.99969 11.0694 0.8546
Without Active Trajectories 0.9815 2.2848 1.5115 12.7503 0.99983 6.6159 0.9193
Without Prior Trajectories 1.0335 2.3063 1.5187 13.0558 0.99983 7.5159 0.9185
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Table 9: Model Performance for Q3 (July–September)
[…]
July Full Context 0.9779 2.0390 1.4279 11.8794 0.99983 7.3058 0.9178
Without Textual Information 1.0305 2.1855 1.4783 12.6284 0.99982 7.5088 0.9118
Without Flight Information 0.9766 2.0451 1.4301 12.1323 0.99983 7.3660 0.9175
Without Weather Information 0.9495 2.0651 1.4371 11.9360 0.99983 6.6536 0.9167
Without METAR 0.9866 2.1627 1.4706 12.1123 0.99982 7.2214 0.9128
Without TAF 1.0063 2.2215 1.4905 12.3320 0.99982 7.4820 0.9104
Without NOTAMs 0.9151 1.9631 1.4011 11.5509 0.99984 6.8400 0.9208
Without All Trajectories 2.9250 18.8846 4.3456 25.9644 0.99846 21.7617 0.2382
Without Focusing Trajectory 1.2792 3.6322 1.9058 13.9341 0.99970 11.2865 0.8535
Without Active Trajectories 0.9996 2.2389 1.4963 12.2851 0.99982 7.9943 0.9097
Without Prior Trajectories 0.9825 2.0861 1.4443 12.2267 0.99983 6.9645 0.9159
August Full Context 0.9675 2.2742 1.5080 12.1986 0.99979 6.5976 0.9183
Without Textual Information 1.0189 2.5999 1.6124 12.6721 0.99976 7.1502 0.9066
Without Flight Information 1.0427 2.4984 1.5806 13.0385 0.99977 8.1249 0.9103
Without Weather Information 1.0539 2.7081 1.6456 12.9250 0.99975 7.0680 0.9028
Without METAR 1.0057 2.6054 1.6141 12.5688 0.99976 6.8738 0.9065
Without TAF 0.9486 2.3424 1.5305 12.0928 0.99978 6.3984 0.9159
Without NOTAMs 0.9755 2.3303 1.5265 12.2459 0.99978 6.6143 0.9163
Without All Trajectories 2.6377 14.8882 3.8585 23.8236 0.99861 20.3841 0.4654
Without Focusing Trajectory 1.4391 5.0106 2.2384 15.7713 0.99953 11.8805 0.8201
Without Active Trajectories 1.0491 2.8388 1.6849 12.6682 0.99974 6.9338 0.8981
Without Prior Trajectories 1.0348 2.6805 1.6372 12.8909 0.99975 7.9558 0.9038
September Full Context 0.8721 1.8700 1.3675 10.3723 0.99982 6.8206 0.9340
Without Textual Information 0.9505 2.1839 1.4778 11.1034 0.99979 7.8916 0.9230
Without Flight Information 0.8799 1.8970 1.3773 10.3290 0.99982 7.3468 0.9331
Without Weather Information 0.9774 2.2235 1.4911 11.2718 0.99979 7.8802 0.9216
Without METAR 0.9334 2.0244 1.4228 11.1036 0.99981 7.6207 0.9286
Without TAF 0.8734 1.9449 1.3946 10.0869 0.99982 7.0707 0.9314
Without NOTAMs 1.0789 2.4187 1.5552 11.9766 0.99977 9.7979 0.9147
Without All Trajectories 2.5926 14.4186 3.7972 23.0578 0.99863 23.6338 0.4912
Without Focusing Trajectory 1.5898 5.7639 2.4008 15.9623 0.99945 15.4688 0.7966
Without Active Trajectories 0.9780 2.2288 1.4929 11.0431 0.99979 8.0546 0.9214
Without Prior Trajectories 0.8919 2.0604 1.4354 10.2853 0.99980 7.1557 0.9273
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
Table 10: Model Performance for Q4 (October-December)
[…]
October Full Context 0.8931 1.8767 1.3699 11.5245 0.99969 7.1497 0.9308
Without Textual Information 1.0133 2.5349 1.5921 12.5664 0.99958 8.0088 0.9065
Without Flight Information 0.8415 1.7910 1.3383 11.0727 0.99970 6.9427 0.9340
Without Weather Information 1.0511 3.0188 1.7375 12.9822 0.99950 8.6614 0.8887
Without METAR 0.8497 1.6838 1.2976 11.0734 0.99972 6.7812 0.9379
Without TAF 0.8852 1.8401 1.3565 11.5888 0.99970 7.1022 0.9321
Without NOTAMs 0.8630 1.8330 1.3539 11.4442 0.99970 7.1753 0.9324
Without All Trajectories 2.3656 13.6659 3.6967 22.8737 0.99774 21.1300 0.4961
Without Focusing Trajectory 1.3790 5.2925 2.3005 15.7475 0.99913 13.4323 0.8049
Without Active Trajectories 0.8553 1.8187 1.3486 11.3867 0.99970 6.5662 0.9329
Without Prior Trajectories 1.1746 4.0501 2.0125 14.1528 0.99933 9.3349 0.8507
November Full Context 0.8932 1.7821 1.3350 11.4918 0.99970 7.4177 0.9373
Without Textual Information 0.9791 2.1326 1.4603 12.6642 0.99964 7.7896 0.9250
Without Flight Information 1.2715 3.6552 1.9119 14.4189 0.99939 11.1492 0.8714
Without Weather Information 0.9327 1.8441 1.3580 12.1507 0.99969 8.0972 0.9351
Without METAR 0.8983 1.7716 1.3310 11.6085 0.99970 6.9296 0.9377
Without TAF 0.8672 1.6654 1.2905 11.4647 0.99972 6.9180 0.9414
Without NOTAMs 0.8392 1.6595 1.2882 11.1810 0.99972 7.1453 0.9416
Without All Trajectories 2.8065 17.6258 4.1983 27.3398 0.99706 23.3370 0.3800
Without Focusing Trajectory 1.1871 3.4078 1.8460 14.4316 0.99943 11.9716 0.8801
Without Active Trajectories 0.9250 1.8833 1.3723 11.9705 0.99969 8.7367 0.9338
Without Prior Trajectories 0.8404 1.6763 1.2947 11.1934 0.99972 6.6889 0.9410
December Full Context 0.8934 1.6342 1.2784 12.9921 0.99983 6.4888 0.9378
Without Textual Information 0.9643 1.9332 1.3904 13.1579 0.99980 7.2932 0.9264
Without Flight Information 0.8881 1.6077 1.2679 12.4841 0.99984 6.9389 0.9388
Without Weather Information 0.9299 1.7443 1.3207 13.1765 0.99982 6.5655 0.9336
Without METAR 0.8770 1.6226 1.2738 12.5185 0.99983 6.5165 0.9382
Without TAF 0.9192 1.6872 1.2989 13.0861 0.99983 6.4581 0.9358
Without NOTAMs 0.7564 1.1624 1.0781 11.3962 0.99988 5.7425 0.9558
Without All Trajectories 2.7659 19.1522 4.3763 26.3006 0.99804 21.0256 0.2711
Without Focusing Trajectory 1.3692 4.1613 2.0399 16.7226 0.99957 12.2092 0.8416
Without Active Trajectories 1.3759 3.9133 1.9782 17.3345 0.99960 9.9423 0.8511
Without Prior Trajectories 1.3759 3.9133 1.9782 17.3345 0.99960 9.9423 0.8511
$\downarrow$ Lower is better, $\uparrow$ Higher is better.
[…]